# XOXO: Stealthy Cross-Origin Context Poisoning Attacks against AI Coding Assistants

### v1

@@ Abstract
AI coding assistants are widely used for tasks like code generation, bug detection, and comprehension. These tools now require large and complex contexts, automatically sourced from various origins—across files, projects, and contributors—forming part of the prompt fed to underlying LLMs. This automatic context-gathering introduces new vulnerabilities, allowing attackers to subtly poison input to compromise the assistant’s outputs, potentially generating vulnerable code, overlooking flaws, or introducing critical errors. We propose a novel attack, Cross-Origin Context Poisoning (XOXO), that is particularly challenging to detect as it relies on adversarial code modifications that are semantically equivalent. Traditional program analysis techniques struggle to identify these correlations since the semantics of the code remain correct, making it appear legitimate. This allows attackers to manipulate code assistants into producing incorrect outputs, including vulnerabilities or backdoors, while shifting the blame to the victim developer or tester. We introduce a novel, task-agnostic black-box attack algorithm GCGS that systematically searches the transformation space using a Cayley Graph, achieving an 83.09% attack success rate on average across five tasks and eleven models, including GPT-4o and Claude 3.5 Sonnet v2 used by many popular AI coding assistants. Furthermore, existing defenses, including adversarial fine-tuning, are ineffective against our attack, underscoring the need for new security measures in LLM-powered coding tools.
@@ 1 Introduction
Figure 1: Comparison between a benign and vulnerable workflow for a developer using GitHub Copilot in a Python-based Django web application project. (a) In the benign workflow, a developer requests a completion for the function SearchQuestionView, and GitHub Copilot generates secure code based on context it gathered for this task. (b) In the vulnerable workflow, an attacker performs Cross-Origin Context Poisoning with a semantics-preserving transformation. As a result, the same code completion request leads to the generation of SQL injection-vulnerable code by GitHub Copilot.
Advances in Large Language Models (LLMs) for code generation and reasoning are revolutionizing software development and maintenance. AI coding assistants integrated into major IDEs like Visual Studio Code [1], JetBrains [2], and Replit [3] have become essential for tasks such as code completion, repair, and bug detection. A 2024 Stack Overflow survey [4] found that 76% of developers are using or planning to adopt AI tools, and Google reports that over 25% of their new code is generated by AI [5].
@@
A key requirement of all AI coding assistants is that they should support real-world software development, extending beyond isolated tasks on curated datasets where LLMs are often tested. These assistants should communicate with LLMs by sending detailed prompts augmented with relevant project context, such as open files, cross-project resources, and prior code. The LLM-generated code should integrate seamlessly with the larger project, adhering to established standards, dependencies, and external APIs. To deliver such context-aware results, coding assistants should ensure that the augmented prompts provide the necessary information for the LLM’s output to align with broader project requirements and conventions. Until recently, achieving this was difficult because earlier transformer models were constrained by small, fixed-size context windows [6], severely limiting prompt size.
@@
However, recent advances in LLMs have significantly accelerated prompt augmentation in coding assistants and other applications, enabling them to leverage larger and more diverse prompts than ever before. Improvements in attention mechanisms and positional encodings now enable modern LLMs to process prompts spanning millions of tokens [7, 8, 9, 10]. As prompts are becoming increasingly larger, making the user manually supply all necessary context becomes exceedingly tedious and impractical.
Therefore, modern LLM-based coding assistants are streamlining development by automatically adding relevant context to LLM prompts. The context is drawn from various sources, such as related files in the project, external libraries, and API usage from other projects. Deciding which content is relevant for a specific task relies on assistant-specific heuristics, such as analyzing task history, identifying co-accessed files, and other assistant-defined criteria. These combined efforts enable LLMs to generate code that aligns with existing specifications and integrates seamlessly into the project [11]
@@
Automatically gathering context from various sources – such as files, projects, and contributions from other developers with varying trust levels – introduces a new attack vector in coding assistants powered by LLMs. These assistants pass the gathered context to LLMs as a unified task request, without differentiating the origin or trustworthiness of each component. Lack of such information makes it difficult to identify whether parts of the context come from potentially untrusted sources, such as files or projects contributed or edited by different developers. Our survey of seven coding assistants (Table I) reveals that all employ automatic context-gathering heuristics, often without developer awareness [11]. Additionally, none provide mechanisms to view, limit, or log the gathered context or associate it with specific queries, primarily due to performance and storage constraints.
@@
In this paper, we introduce Cross-Origin Context Poisoning (XOXO), a practical, stealthy, and hard-to-trace attack that exploits the attack vector of automatically gathered context in AI coding assistants. Unlike prompt injection attacks, which directly alter the prompt’s meaning by changing its semantics, XOXO attack manipulates the assistant’s context through subtle modifications that preserve both the natural language semantics of the prompt and the functional behavior of the code. XOXO attack achieves this by introducing minor changes—such as renaming variables or adding harmless code snippets—to contributed files. These subtle modifications preserve the prompt’s meaning and the code’s functionality while aiming to mislead the LLM into producing insecure or faulty outputs, as shown in Figure 2. Specifically, when a malicious developer collaborating on the same project or a related project introduces these changes, the AI assistant automatically includes them in its context by leveraging its heuristics without the victim developer’s knowledge. This approach makes the attack both covert and difficult to trace. As a result, the victim developer might unknowingly integrate compromised code, believing it to be secure as AI assistants have been shown to make developers over-confident [12].
@@
The lack of transparency, origin isolation, and traceability of prompt components in current coding assistant architectures significantly amplifies the stealth and impact of XOXO. While malicious developers can manually insert vulnerable code without AI, AI coding assistants allow them to obscure their actions, shifting responsibility to the victim developer. This lack of accountability makes tracing the origin of vulnerabilities exceptionally difficult. Figure 1 illustrates a real-world attack where a seemingly innocuous, semantically identical variable renaming (e.g., changing USE_RAW_QUERIES to RAW_QUERIES) in GitHub Copilot’s automatically gathered context bypasses its AI-powered vulnerability prevention system [13], resulting in the generation of vulnerable code with an exploitable SQL injection flaw [^1] We responsibly disclosed this attack to the GitHub Copilot team..
@@
We present Greedy Cayley Graph Search (GCGS), an efficient black-box attack algorithm that identifies semantics-preserving adversarial transformations causing Large Language Models (LLMs) to generate buggy and even vulnerable code, or to fail on reasoning tasks such as defect and clone detection. A key challenge in this task is the vast transformation space, as the number of possible semantics-preserving transformations is combinatorially large. Our approach begins with a set of generator transforms—basic operations that preserve program semantics, such as variable renaming—and composes them to form a free group. This composition inherently preserves semantics and allows us to systematically create and analyze all possible combinations of these transformations. We represent this structure using a Cayley Graph, which in the context of a free group forms a tree where each path corresponds to a unique sequence of generator transforms. GCGS leverages the monotonicity property—empirically observed to hold on average for the LLMs we tested—where composing additional transforms either increases or maintains the misprediction rate. By employing a greedy traversal strategy on the Cayley graph, GCGS efficiently identifies and composes atomic transformations that most significantly degrade the model’s confidence, generating adversarial inputs that cause LLMs to mispredict.
@@
Since adversarial transforms depend on the underlying LLMs, we evaluate GCGS directly on these foundation models. To assess our approach, we conducted extensive experiments on code generation and code reasoning tasks. First, we tested GCGS’s ability to deceive state-of-the-art LLMs into generating buggy code on two popular benchmarks, HumanEval+ and MBPP+, successfully targeting top performers such as GPT-4o, Claude 3.5 Sonnet, and Llama 3 [14]. Our attack achieved average success rates of 86.47% attacking SoTA models on HumanEval+ and 94.83% on MBPP+, requiring only 156 and 63 queries on average, respectively, and maintained high naturalness with CodeBLEU scores above 96.01. Second, we evaluated GCGS’s ability to make LLMs generate vulnerable code using CWEval [15], a benchmark designed to test LLMs’ secure code generation capabilities. Our approach made SoTA LLMs generate vulnerable yet fully functional code, with attack success rates up to 75%. Finally, we evaluated code reasoning capabilities through two CodeXGLUE tasks: Defect Detection and Clone Detection. Against fine-tuned models like CodeBERT, GraphCodeBERT, and CodeT5+, our approach outperformed the best-performing existing attacks by up to 23.20 percentage points for defect detection and 38.28 percentage points for clone detection while using 65.14% and 43.64% queries, respectively, and increasing code similarity. These results reveal the broad vulnerability of current coding LLMs (and the AI coding assistants using them) to semantics-preserving transformations and underscore the need for robust defenses.
@@
  • Identifying Cross-Origin Context Poisoning (XOXO), a practical and highly stealthy attack vector in AI coding assistants that exploits automatic and extensive context inclusion without transparency.
  • Demonstrating an end-to-end practical attack how subtle context changes, such as identifier replacements, can manipulate AI coding assistants like GitHub Copilot to generate vulnerable code.
  • Introducing the Greedy Cayley Graph Search algorithm, which efficiently identifies semantics-preserving transformations that deceive models into generating buggy and vulnerable code and fail on reasoning tasks. We are working on making the source code public.
  • Performing extensive evaluation and showing that GCGS can achieve an average 90.65% success rate on in-context code generation against SoTA models like GPT-4o, Claude 3.5 Sonnet, and Llama 3, injecting vulnerabilities with success rates up to 75%, as well as outperforming previous attacks on code reasoning tasks with up to 38.28 percentage points higher success rates, 43.64% fewer queries, and increased code similarity.
@@ 2 AI Assistant Background
TABLE I: Survey of AI coding assistants detailing context origins, tasks supported, and if the backend model the assistants query is configurable. Different context pulling methods are–Intra-File: context pulled from the same file, Inter-File: context pulled across multiple files, Inter-Project: context pulled across multiple projects.
@@
Software tasks supported by these assistants fall into two broad categories: code generation (e.g., code completion, refactoring, or repair) and code reasoning (e.g., defect identification, code summarization, locating references to publicly available code to avoid licensing infringements).
AI Coding Assistant Interfaces. Developers interact with AI coding assistants through in-line code or chat interfaces. Code generation tasks can utilize either interface, while code reasoning primarily employs chat. Developers can formulate their query using code and/or natural language, with the assistant generating appropriate prompts for the underlying LLM and filling responses at the user-specified locations.
Embedded within IDEs, AI coding assistants leverage the IDEs’ project structure and metadata to reason over codebases effectively. A project, comprising multiple source files and associated resources, is managed within a workspace—an IDE environment that encapsulates all project-related files, configuration settings, build scripts, and dependencies pertinent to that project.
Context Gathering. AI coding assistants handle tasks differently based on the interface. In in-line suggestion mode, these assistants eagerly query the LLM with each developer-triggered event (e.g., new code or comments), often maintaining a session cache for repeated, identical queries, while chat interfaces operate in a request-response format. In both cases, the assistant augments prompts with relevant context to improve the LLM’s suggestions.
To understand the AI coding assistants’ context collection and prompt augmentation mechanisms, we intercepted network communication between these assistants and their underlying LLMs as we tried to complete various coding tasks. Details about traffic interception are provided in § B.1. Through this analysis, we identified three primary context collection methods (detailed in Table I):
  • Inter-project: Accessing code from projects other than the one containing the file being modified
  • Inter-file: Fetching context from files within the current project, excluding the file being modified
  • Intra-file: Fetching context from within the file being modified
The assistants always include Intra-file context, regardless of the task. Inter-file and Inter-project context is gathered primarily through simultaneously open files, chat interface history, and dependency declarations (e.g., include in C, import in Python).
Beyond these automatic mechanisms, these assistants support additional context gathering approaches. When explaining code snippets, they may incorporate code from files referenced through function calls, class inheritance, or similar dependencies. These references are retrieved using the IDE’s native code referencing mechanisms. Developers can also explicitly specify additional files to include in the context. These explicit interfaces cannot be used to exclude any files from the automatically gathered context. Details about these explicit context interfaces are provided in § B.2.
@@ 3 Threat Model
In our threat model, we consider a malicious developer with commit privileges to a collaborative project with the victim developer. This attacker can introduce semantics-preserving adversarial transformations to source code files—modifications that alter the code’s structure while maintaining its functional behavior. These transformations are designed to deceive AI coding assistants by altering the code in ways that will fool LLMs while preserving the intended semantics. These changes evade both manual code review, by appearing natural to human reviewers, and automated regression testing or other automated analysis tools, due to their semantic equivalence. The attacker is assumed to know which parts of the codebase the victim developer is actively working on, as well as which AI assistant or LLM the victim is using, a reasonable assumption given that such information can often be inferred from issue trackers, feature requests, or past project commit history. Furthermore, the attacker can determine which files or parts of the project’s codebase will be fetched by an AI coding assistant for the target section, as the procedure to reverse-engineer the AI coding assistant’s behavior shown in § B.1 can be replicated. Once the attacker’s modified code is merged into the main branch, it will be pulled in by the victim developer through normal version control operations. As shown in Table I and discussed previously, AI assistants utilize multiple avenues to gather context for their code development suggestions. Consequently, the attacker’s modifications become part of this context, leading the AI assistants to generate vulnerable output based on the task being performed.
@@ 4 An End-to-End XOXO Attack
To demonstrate XOXO attack’s severity, we present a real-world end-to-end attack in a practical setting, illustrated in Figure 1. We target GitHub Copilot integrated into VSCode—a widely adopted AI coding assistant with extensive security safeguards [23]—making it an ideal candidate for validating XOXO attack’s effectiveness. Our attack simulation takes place in a Python web application using the Django Framework [24], where we demonstrate how a malicious developer with commit privileges can successfully leverage XOXO attack to manipulate GitHub Copilot into generating code containing an SQL injection vulnerability.
The victim developer is working on a web-based polls application, implementing a feature request for an endpoint where users can search questions using a question_text parameter. Given Django’s model-view-controller architecture, the attacker correctly anticipates that the developer will implement this feature in views.py. Knowing that GitHub Copilot automatically incorporates context from the entire file, the attacker commits a subtle, semantics-preserving transformation upstream by renaming a global variable from USE_RAW_QUERIES to RAW_QUERIES. Through prior experimentation with the model, the attacker has determined that this change triggers GitHub Copilot to generate code that uses unsanitized user-supplied input in SQL queries, as shown on the right in Figure 1, whereas it previously suggested secure versions using Django’s built-in input sanitization, as shown on the left in Figure 1. The figure illustrates how this seemingly benign change, once merged into the main branch and pulled by the victim developer, manipulates GitHub Copilot into generating vulnerable code instead of the secure implementation it would normally suggest.
To validate the attack’s effectiveness and establish its reliability, we conducted extensive testing across multiple Copilot sessions. The assistant consistently generated vulnerable code, demonstrating the stability predicted by our threat model. This consistency is attributed to Copilot’s low temperature setting (0.1), which ensures minimal non-determinism in the generated outputs. Through systematic comparison of generations with and without the transformation, we confirmed that the vulnerability appears only when the context is poisoned, definitively establishing XOXO attack as the root cause of the vulnerable output. Furthermore, we successfully triggered this vulnerable suggestion even when moving the variable RAW_QUERIES to models.py and merely importing it into the target views.py file, demonstrating the attack’s resilience across file boundaries. We verified the functionality of this XOXO attack instance on the latest GitHub Copilot version available at the time of submission and responsibly disclosed the vulnerability to the vendor.

### v2

@@ Abstract
AI coding assistants are widely used for tasks like code generation. These tools now require large and complex contexts, automatically sourced from various origins—across files, projects, and contributors—forming part of the prompt fed to underlying LLMs. This automatic context-gathering introduces new vulnerabilities, allowing attackers to subtly poison input to compromise the assistant’s outputs, potentially generating vulnerable code or introducing critical errors. We propose a novel attack, Cross-Origin Context Poisoning (XOXO), that is challenging to detect as it relies on adversarial code modifications that are semantically equivalent. Traditional program analysis techniques struggle to identify these perturbations since the semantics of the code remains correct, making it appear legitimate. This allows attackers to manipulate coding assistants into producing incorrect outputs, while shifting the blame to the victim developer. We introduce a novel, task-agnostic, black-box attack algorithm GCGS that systematically searches the transformation space using a Cayley Graph, achieving a 75.72% attack success rate on average across five tasks and eleven models, including GPT 4.1 and Claude 3.5 Sonnet v2 used by popular AI coding assistants. Furthermore, defenses like adversarial fine-tuning are ineffective against our attack, underscoring the need for new security measures in LLM-powered coding tools.
@@ 1 Introduction
Figure 1: Comparison between a benign and vulnerable workflow for a developer using GitHub Copilot in a Python-based Django web application project. (a) In the benign workflow, a developer requests a completion for the class SearchQuestionView, and GitHub Copilot generates secure code based on context it gathered for this task. (b) In the vulnerable workflow, an attacker performs Cross-Origin Context Poisoning with a semantics-preserving transformation. As a result, the same code completion request makes GitHub Copilot generate SQL injection-vulnerable code.
@@
AI coding assistants integrated into major IDEs like Visual Studio Code [24] have become essential for tasks such as code completion and repair. A 2024 Stack Overflow survey [45] found that 76% of developers are using or planning to adopt AI tools. A key requirement of all AI coding assistants is supporting real-world software development. These assistants should communicate with Large Language Models (LLMs) by sending detailed prompts augmented with relevant project context to ensure LLM-generated code integrates seamlessly with the larger project. Delivering such results requires the coding assistants to augment the prompts with relevant information from the project. Until recently, achieving this was difficult because earlier transformer models were constrained by small context windows [53], severely limiting prompt size.
@@
However, recent advances in LLMs have significantly accelerated prompt augmentation in coding assistants and other applications, enabling them to leverage larger and more diverse prompts than ever before [16, 41, 11, 50]. As prompts are becoming increasingly larger, making the user manually supply all necessary context becomes exceedingly tedious and impractical. Consequently, modern LLM-based coding assistants are streamlining development by automatically adding relevant context to LLM prompts. Assistants draw context from multiple sources such as project-related files and external library usage, which helps LLMs generate code that aligns with existing specifications [54].
@@
Automatically gathering context from various sources – often including contributions from other developers with varying trust levels – introduces a new attack vector in coding assistants. These assistants pass the gathered context to LLMs as a unified task request, without differentiating the origin or trustworthiness of each component. Lack of such information makes it difficult to identify whether parts of the context come from potentially untrusted sources. Our survey of seven coding assistants (Table 14) reveals that all employ automatic context-gathering heuristics, often without developer awareness [54]. Additionally, none provide mechanisms to view, limit, or log the gathered context or associate it with specific queries, primarily due to performance and storage constraints.
@@
In this paper, we introduce Cross-Origin Context Poisoning (XOXO), a practical, stealthy, and hard-to-trace attack targeting automatically gathered context in AI coding assistants. Unlike prompt injection attacks, which directly alter prompts, XOXO attack manipulates the assistant’s context through subtle modifications that preserve both the prompt’s semantics and the code’s functional behavior. XOXO attack achieves this by introducing minor changes to contributed files, such as renaming variables or adding dead code snippets. These subtle modifications aim to mislead the LLM into producing insecure or faulty outputs, as shown in Figure 2. Specifically, when a malicious developer collaborating on the same project or a related project introduces these changes, the AI assistant automatically includes them in its context without the victim developer’s knowledge, making the LLM generate faulty or even vulnerable code. This approach makes the attack both covert and difficult to trace. Consequently, the victim developer might unknowingly integrate compromised code, falsely trusting AI assistant-generated code as secure [47].
@@
The lack of transparency and traceability of prompt components in coding assistant architectures significantly amplifies the stealth and impact of XOXO. While malicious developers can manually insert vulnerable code, AI coding assistants allow them to obscure their actions, shifting responsibility to the victim developer. This lack of accountability makes tracing the origin of vulnerabilities exceptionally difficult. Figure 1 illustrates a real-world attack where a seemingly innocuous, semantically identical variable renaming (e.g., changing USE_RAW_QUERIES to RAW_QUERIES) in GitHub Copilot’s automatically gathered context bypasses its AI-powered vulnerability prevention system [72], resulting in the generation of vulnerable code with an exploitable SQL injection flaw [^1] We responsibly disclosed this attack to the GitHub Copilot team..
@@
Moreover, we present Greedy Cayley Graph Search (GCGS), an efficient black-box attack algorithm that identifies semantics-preserving adversarial transformations causing LLMs to generate buggy or vulnerable code or to fail on defect and clone detection. A key challenge in this task is the vast transformation space, as the number of possible semantics-preserving transformations is combinatorially large. Our approach begins with a set of generator transforms – basic operations that preserve program semantics, such as variable renaming – and composes them to form a free group. This composition preserves semantics and allows us to systematically explore promising combinations of these transformations. We represent this structure using a Cayley Graph, which in the context of a free group forms a tree where each path corresponds to a unique sequence of generator transforms. GCGS leverages the monotonicity property where composing additional transforms either increases or maintains the misprediction rate. By employing a greedy traversal strategy on the Cayley graph, GCGS efficiently identifies and composes atomic transformations that significantly degrade the model’s confidence, generating adversarial inputs that cause LLMs to mispredict.
@@
To assess the efficacy of our approach, we conduct experiments on code generation and reasoning tasks. First, GCGS can deceive state-of-the-art LLMs (including GPT 4.1, Claude 3.5 Sonnet v2, and Qwen 2.5 Coder 32B) into generating buggy code with average attack success rates (ASR) of 83.67%. Second, on CWEval [46], a benchmark testing LLM’s secure code generation capabilities, GCGS can make SoTA LLMs generate vulnerable yet fully functional code, with ASR up to 66.67%. Finally, on two code reasoning tasks (Defect Detection and Clone Detection), GCGS outperforms existing best-performing attacks against fine-tuned models with ASR improvements up to 38.28 percentage points.
@@
  • Identifying Cross-Origin Context Poisoning (XOXO), a practical and highly stealthy attack vector in AI coding assistants that exploits automatic, mixed-origin context inclusion.
  • Demonstrating an end-to-end real-world attack that manipulates AI coding assistants like GitHub Copilot into generating vulnerable code using subtle context changes (e.g., variable renaming).
  • We introduce Greedy Cayley Graph Search (GCGS), an efficient algorithm for finding semantics-preserving transformations that cause models to fail on code generation and reasoning tasks.
  • Showcasing through extensive evaluation that GCGS can achieve an average 83.67% attack success rate on in-context code generation against SoTA models like GPT 4.1, Claude 3.5 Sonnet v2, and Qwen 2.5 Coder 32B, injecting vulnerabilities with success rates up to 66.67%.
@@ 2 AI Assistant Background
Software tasks supported by these assistants fall into two broad categories: code generation (e.g., code completion, refactoring) and code reasoning (e.g., defect identification, code summarization).
AI Coding Assistant Interfaces. Developers interact with AI coding assistants through in-line code or chat interfaces. Developers can formulate their query using code and/or natural language, with the assistant generating appropriate prompts for the underlying LLM and filling responses at the user-specified locations.
Embedded within IDEs, AI coding assistants leverage the IDEs’ project structure and metadata to reason over the codebase. A project, comprising multiple source files, is managed within a workspace – an IDE environment that encapsulates all source code files and configuration scripts.
Context Gathering. In in-line suggestion mode, these assistants eagerly query the LLM with each developer-triggered event (e.g., new code or comments), while chat interfaces operate in a request-response format. In both cases, the assistant augments prompts with relevant context to improve the LLM’s suggestions. To understand the AI coding assistants’ context collection, we intercepted their network communication while completing various coding tasks (details provided in § E.1). Through this analysis, we identified three primary context collection methods (detailed in Table 14): (i) Inter-project: Accessing code from projects other than the one containing the file being modified, (ii) Inter-file: Fetching context from files within the current project, excluding the file being modified, and (iii) Intra-file: Fetching context from within the file being modified.
The assistants always include intra-file context. Inter-file and inter-project context is gathered primarily through simultaneously open files and chat interface history. Beyond these automatic mechanisms, these assistants support additional developer-driven context gathering approaches (see § E.2).
@@ 3 Threat Model
In our threat model, we consider a malicious developer with commit privileges to a collaborative project with the victim developer. This attacker can introduce semantics-preserving transformations to source code files – modifications that alter the code’s structure while maintaining its functional behavior. These changes evade manual code review by appearing natural to human reviewers and automated regression testing due to their semantic equivalence. The attacker is assumed to know which parts of the codebase the victim developer is working on and which AI assistant or LLM the victim is using, a reasonable assumption given that such information can often be inferred from issue trackers or feature requests. Furthermore, the attacker can determine which parts of the project’s codebase will be fetched by an AI assistant for the target section by reverse-engineering AI assistant’s behavior (§ E.1). Once the attacker’s modified code is merged into the main branch, it will be pulled in by the victim developer through normal version control operations. Since AI assistants utilize multiple avenues to gather context for their code development suggestions (§ 2), the attacker’s modifications become part of this context, leading them to generate vulnerable output.
@@ 4 An End-to-End XOXO Attack
To demonstrate XOXO attack’s severity, we present a real-world end-to-end attack in a practical setting (Figure 1). We target GitHub Copilot integrated into VSCode – a popular AI coding assistant with extensive security safeguards [23]. In a Python web application using the Django Framework [18], we show how a malicious developer with commit privileges can leverage the XOXO attack to manipulate Copilot into generating code containing an SQL injection vulnerability.
The victim developer is developing a web-based polls application, implementing a feature request for an endpoint to search questions using a question_text parameter. Given Django’s architecture, the attacker correctly anticipates that the developer will implement this feature in views.py. Knowing that Copilot automatically incorporates context from the entire file, the attacker commits a subtle, semantics-preserving transformation by renaming a global variable from USE_RAW_QUERIES to RAW_QUERIES. Through prior experimentation, the attacker knows this change triggers Copilot to generate code that uses unsanitized user-supplied input in SQL queries (shown on the right in Figure 1), whereas it previously suggested secure versions using Django’s input sanitization (shown on the left in Figure 1). The figure illustrates how this benign change, once merged into the main branch and pulled by the victim developer, manipulates Copilot into generating vulnerable code.
To validate the attack’s effectiveness and reliability, we tested it across multiple Copilot sessions. The assistant consistently generated vulnerable code which can be attributed to Copilot’s low temperature setting (0.1). The low temperature ensures minimal non-determinism in the generated outputs. Through systematic comparison of generations with and without the transformation, we confirmed that the vulnerability appears only when the context is poisoned, establishing XOXO attack as the root cause of the vulnerable output. Furthermore, we successfully triggered this vulnerable suggestion even when moving the variable RAW_QUERIES to models.py and merely importing it into the target views.py file, demonstrating the attack’s resilience across file boundaries. We verified the functionality of this XOXO attack instance on Copilot versions 1.239-1.243 and responsibly disclosed the vulnerability to the vendor, who addressed it by the time of this submission.